# NAMI: Efficient Image Generation via Progressive Rectified Flow Transformers

### v1

@@ Abstract
Flow-based transformer models for image generation have achieved state-of-the-art performance with larger model parameters, but their inference deployment cost remains high. To enhance inference performance while maintaining generation quality, we propose progressive rectified flow transformers. We divide the rectified flow into different stages according to resolution, using fewer transformer layers at the low-resolution stages to generate image layouts and concept contours, and progressively adding more layers as the resolution increases. Experiments demonstrate that our approach achieves fast convergence and reduces inference time while ensuring generation quality. The main contributions of this paper are summarized as follows: (1) We introduce progressive rectified flow transformers that enable multi-resolution training, accelerating model convergence; (2) NAMI leverages piecewise flow and spatial cascading of Diffusion Transformer (DiT) to rapidly generate images, reducing inference time by 40% to generate a 1024×1024 resolution image; (3) We propose NAMI-1K benchmark to evaluate human preference performance, aiming to mitigate distributional bias and prevent data leakage from open-source benchmarks. The results show that our model is competitive with state-of-the-art models.
@@ 1 Introduction
Over the past year, text-to-image (T2I) models based on both diffusion [45, 17, 39, 10] and autoregressive approaches [3, 8, 36, 38] have achieved significant advancements in image generation quality and inference efficiency. The diffusion model approach, represented by SD3 [11] and FLUX [22], leverages rectified flow [28, 1, 25] and MM-DiT [11] architectures to achieve excellent performance. However, the increased model parameters have significantly raised training and inference costs, making commercialization more challenging. Concurrently, autoregressive (AR) models generate images through next-token prediction [35, 31]. These methods use vector quantization variational autoencoder (VQVAE) technology [49] to discretize image tokens, and then predict the next component of the image based on the previously generated tokens. The size of the codebook plays a critical role in determining both the efficiency of image generation and the image quality. Currently, AR models such as LlamaGen [47] and VAR [48] exhibit significant advantages in inference speed, but their image generation quality still falls short compared to diffusion models.
@@
In this paper, we propose a spatiotemporal separation progressive framework, NAMI. We divide the rectified flow into different stages according to resolution, using fewer transformer layers at the low-resolution stages to generate image layouts and concept contours, and progressively adding more layers as the resolution increases. As shown in the Figure 2, we visualize the image generation process of FLUX-dev [22] and NAMI-2B under the same inference steps. Despite using a lower resolution and fewer model parameters in the early stages, NAMI-2B still effectively generates image layouts and concepts. Furthermore, considering the overall process, NAMI-2B achieves a faster generation speed.
Current benchmarks such as GenEval [12], DPG-Benchmark [19], and ImageReward [53], are primarily used to evaluate a model’s performance. However, these benchmarks suffer from issues such as limited caption quantity, distribution biases, and data leakage. Therefore, a more comprehensive benchmark is essential for accurately evaluating a model’s capabilities. We combine open benchmarks, AI-generated, and human-created test prompts to construct a multidimensional evaluation dataset named NAMI-1K.
@@
    The NAMI-2B employs temporal separation and spatial cascading within the model to efficiently generate images. Using the same model size, our approach reduces the time required to generate 1024 resolution images by 40%, while still maintaining a high level of image quality. Furthermore, as resolution increases, we observe even greater performance improvements. As shown in Figure 1.
@@ 2 Related Work
Diffusion models [45, 17, 39, 10] have become a powerful framework in image generation, with latent diffusion models further enhancing both quality and efficiency. DiT [32] has introduced a transformative change by replacing the conventional UNet [40, 42, 9] architecture with a transformer [50]-based framework. This advancement has led to the development of models such as PixArt-$\alpha$ [5], Hunyuan-DiT [24], SD3 [11], and FLUX [22], which efficiently integrate multimodal information to improve text-to-image generation, thereby demonstrating the effectiveness of transformer-based approaches. Notably, SD3 and FLUX employ rectified flow [28, 1, 25], which connects data and noise on a straight line, improving both training and inference efficiency. Additionally, the MM-DiT [11] block has been proposed, achieving state-of-the-art generation performance. Our work builds upon MM-DiT as the fundamental module while adopting rectified flow as a specific forward path selection strategy.
@@
Figure 3: Overview of NAMI: The left figure shows the progressive flow transformers of NAMI, where the same color represents the same module. The right figure illustrates the method of obtaining the endpoint for each window. Specifically, we divide the image generation process into $K$ resolution stages and the entire flow is divided into $K$ time windows, where adjacent stages are connected through upsampling and downsampling functions. We use fewer transformer layers at the low-resolution stages to generate image layouts and concept contours, progressively adding more layers as the resolution increases.
@@ 3.1 Preliminary Study
[…]
@@ 3.2 Progressive Rectified Flow Transformers
Our method treats the image generation process as a transformation from a low-resolution image containing coarse-grained concepts to a high-resolution image with rich details. At low resolution, we employ piecewise flow [56] and partial modules to reduce computational costs. As shown in Figure 3 , we divide the image generation into $K$ resolution stages, where $Up(.,.)$ denotes the upsampling function and $Down(.,.)$ denotes the downsampling function. For flow processing, we refer to pyramid flow matching [20] and divide the entire flow into $K$ time windows $\{[t_{k-1},t_{k}]\}_{k=1}^{K}$, where $01=t_{K}>\dots>t_{k}>t_{k-1}>\dots>t_{0}=$ 0. Linear interpolation is also used between noise and data distribution, that is, $x_{t_{k}}=t_{k}x_{0}+(1-t_{k})x_{1}$. The starting point for each time window is given by $\hat{x}_{s_{k}}=Up(Down(x_{t_{k-1}},2^{k+1}))$, which is obtained by upsampling the endpoint of the previous stage. The end point is given by $\hat{x}_{e_{k}}=Down(x_{t_{k}},2^{k})$, which is obtained by downsampling $x_{1}$ and adding noise. The flow within the window is given by:
@@
[…]
@@ 3.3 Multi-Resolution Training
2:  Create $K$ time windows $[t_{k-1},t_{k}]$ for $k=1$ to $K$ with $t_{K}=1,t_{0}=0$
@@
5:     for each time window $[t_{k-1},t_{k}]$ do
@@
$\hat{x}_{s_{k}}=\begin{cases}x_{0}&\text{if }k=1\\ Up(Down(x^{\prime}_{t_{k-1}},2))&\text{otherwise}\end{cases}$
9:        Compute the end point $\hat{x}_{e_{k}}=x^{\prime}_{t_{k}}$
10:        Sample $t^{\prime}\in[t_{k-1},t_{k}]$, compute $\hat{x}_{t}=t^{\prime}\hat{x}_{s_{k}}+(1-t^{\prime})\hat{x}_{e_{k}}$
11:        Compute the loss:
@@
12:     end for
13:     Weighting the losses for different time windows and perform backpropagation
14:  until convergence
Considering knowledge sharing between modules and efficient training, we propose a multi-resolution progressive approach during the training process. Different resolutions of data can be used at different stages, and the optimization follows the approach in Section 3.2. The training process is shown in Algorithm 1. Unlike the conventional method of training at low resolutions first and then fine-tuning at higher resolutions separately, our approach enables the joint training of data with different resolutions simultaneously. This facilitates knowledge sharing within the model and helps prevent catastrophic forgetting during fine-tuning at high resolutions. At each stage, we train using images with resolutions greater than or equal to the current resolution. The $Down(.,.)$ function is applied to downsample the images to the corresponding resolution, from which the starting and ending points are computed to obtain the target and loss for that stage. The loss of different stages is jointly optimized by dynamically adjusting their weights according to the training process. Additionally, during training, we also follow SD3 [11] and use logit-normal sampling for each stage. To enable CFG [16], we apply a 0.1 probability for random dropping of the prompt.
@@ 3.4 Inference
Algorithm 2 Inference with Progressive Rectified Flow Transformers
  Input: Noise $x_{0}\sim N(0,I)$, trained $\theta_{k}=\{m_{1}\oplus\dots\oplus m_{K}\}_{k=1}^{K}$
2:  for $k=1$ to $K$ do
     Use $v_{\theta_{k}}$ and the start point $\hat{x}_{s_{k}}$ for its time window to compute the end point $\hat{x}_{e_{k}}$
4:     Use $\hat{x}_{e_{k}}$ for nearest upsampling and renoising to compute the next stage start point $\hat{x}_{s_{k+1}}$
  end for
6:  Output: Generated sample $x_{1}$
During inference, we sample a noise at the minimum resolution, start from stage $k=1$ and proceed until $k=K$, where at each stage, we continue with the standard sampling algorithms of the rectified flow. For the jump points between stages, we follow pyramid flow matching [20] using nearest upsampling with renoising. Our approach of progressive rectified flow transformers can reduce 40% time during inference. The inference process is shown in Algorithm 2.
@@ 4.2 Experiments Details
We trained the NAMI-2B model using the multi-resolution simultaneous training strategy described in Section 3.3. At the early stage of training, we set the learning rate to 1e-4 and sampled three resolutions within each batch at a ratio of 4:2:1 (from low to high resolution) for 120k steps. Subsequently, we adjusted the sampling ratio to 2:4:1 and continued training for a further 80k steps. Finally, we modified the ratio to 1:2:4, reduced the learning rate to 5e-5, and trained for 50k steps to obtain the final NAMI-2B model.
@@ 4.2.1 Quantitative Comparison
Methods Params (B) Overall Single Two Counting Colors Position CA
FLUX-dev [22] 12.0 0.67 0.99 0.81 0.79 0.74 0.20 0.47
FLUX-schnell [22] 12.0 0.71 0.99 0.92 0.73 0.78 0.28 0.54
LUMINA-Next [57] 2.0 0.46 0.92 0.46 0.48 0.70 0.09 0.13
SDXL [34] 2.6 0.55 0.98 0.74 0.39 0.85 0.15 0.23
Hunyuan-DiT [24] 1.5 0.63 0.97 0.77 0.71 0.88 0.13 0.30
SD3-medium [11] 2.0 0.62 0.98 0.74 0.63 0.67 0.34 0.36
Sana-1.6B [52] 1.6 0.66 0.99 0.77 0.62 0.88 0.21 0.47
NAMI-2B 2.0 0.65 0.99 0.78 0.64 0.82 0.20 0.45
Table 2: Comparison of different methods on GenEval. With highlight the best, second best entries. CA means Color Attribution.
We found that GenEval [12] and DPG-Benchmark [19] primarily focus on text-image alignment, and many models exhibit overfitting on open-source benchmarks.
Methods Params (B) Overall Global Entity Attribute Relation Other
FLUX-dev [22] 12.0 84.0 82.1 89.5 88.7 91.1 89.4
FLUX-schnell [22] 12.0 84.8 91.2 91.3 89.7 86.5 87.0
LUMINA-Next [57] 2.0 74.6 82.8 88.7 86.4 80.5 81.8
SDXL [34] 2.6 74.7 83.3 82.4 80.9 86.8 80.4
Hunyuan-DiT [24] 1.5 78.9 84.6 80.6 88.0 74.4 86.4
PixArt-$\Sigma$ [7] 0.6 80.5 86.9 82.9 88.9 86.6 87.7
SD3-medium [11] 2.0 84.1 87.9 91.0 88.8 80.7 88.7
Sana-1.6B [52] 1.6 84.8 86.0 91.5 88.9 91.9 90.7
NAMI-2B 2.0 83.8 90.3 89.4 88.3 88.0 90.6
Table 3: Comparison of different methods on DPG-Benchmark. With highlight the best, second best entries.
@@ 4.2.2 Human Evaluation
To address the data leakage of open-source benchmarks and to accurately evaluate the model’s performance, We propose the NAMI-1K evaluation dataset from the perspective of human preference performance.
@@
As shown in Figure 5, the prompt lengths in GenEval are primarily concentrated within 10 words, while those in the DPG-benchmark are mainly distributed between 50 and 80 words. In contrast, NAMI-1K exhibits a diverse distribution across different lengths within 120 words. As described in Appendix  A. Detailed Description of NAMI-1K, compared to GenEval and the DPG-benchmark, the topic distribution of NAMI-1K is more comprehensive and balanced.
@@
Evaluation Results The evaluation metrics are presented in Table 4. Although our model still trails behind FLUX-dev [22], which have considerably larger parameter sizes, it shows a clear advantage over SD3-medium [11], Infinity [14], Hunyuan-DiT [24], and SANA [52], which have comparable parameter sizes. As illustrated in Figure 4, a comparison of the generation results is provided between different methods.
@@
Model Params(B) Rele Cohe Aes Real Overall
Flux-dev [22] 12.0 83.93 83.28 84.26 90.16 85.05
SD3-medium [11] 2.0 75.74 65.90 61.64 75.74 69.97
Infinity [14] 2.0 76.39 65.25 61.97 74.43 69.77
Hunyuan-DiT [24] 1.5 74.43 69.51 63.93 64.92 68.95
SANA [52] 1.6 75.41 62.30 60.00 72.46 67.80
NAMI-2B 2.0 76.07 66.89 62.30 76.72 70.69
Table 4: Human evaluation results on NAMI-1K dataset. Rele & Cohe & Aes & Real means: Relevance & Coherence & Aesthetic & Realism. With highlight the best, second best entries.

### v2

# NAMI: Efficient Image Generation via Bridged Progressive Rectified Flow Transformers

@@ Abstract
Flow-based Transformer models have achieved state-of-the-art image generation performance, but often suffer from high inference latency and computational cost due to their large parameter sizes. To improve inference efficiency without compromising quality, we propose Bridged Progressive Rectified Flow Transformers (NAMI), which decompose the generation process across temporal, spatial, and architectural demensions. We divide the rectified flow into different stages according to resolution, and use a BridgeFlow module to connect them. Fewer Transformer layers are used at low-resolution stages to generate image layouts and concept contours, and more layers are progressively added as the resolution increases. Experiments demonstrate that our approach achieves fast convergence and reduces inference time while ensuring generation quality. The main contributions of this paper are summarized as follows: (1) We introduce Bridged Progressive Rectified Flow Transformers that enable multi-resolution training, accelerating model convergence; (2) NAMI leverages piecewise flow and spatial cascading of Diffusion Transformer (DiT) to rapidly generate images, reducing inference time by 64% for generating 1024×1024 resolution images; (3) We propose a BridgeFlow module to align flows between different stages; (4) We propose the NAMI-1K benchmark to evaluate human preference performance, aiming to mitigate distributional bias and comprehensively assess model effectiveness. The results show that our model is competitive with state-of-the-art models.
@@ 1 Introduction
Over the past year, text-to-image (T2I) models based on both diffusion [44, 16, 38, 9] and autoregressive approaches [3, 7, 35, 37] have achieved significant advancements in image generation quality and inference efficiency. The diffusion model approach, represented by SD3 [10] and FLUX [21], leverages rectified flow [27, 1, 24] and MM-DiT [10] architectures to achieve excellent performance. However, the increased model parameters have significantly raised training and inference costs, making commercialization more challenging. Concurrently, autoregressive (AR) models generate images through next-token prediction [34, 30]. These methods use vector quantization variational autoencoder (VQVAE) technology [47] to discretize image tokens, and then predict the next component of the image based on the previously generated tokens. The size of the codebook plays a critical role in determining both the efficiency of image generation and the image quality. Currently, AR models such as LlamaGen [46] and Infinity [13] exhibit significant advantages in inference speed, but their image generation quality still falls short compared to diffusion models.
@@
In this paper, we propose a spatiotemporal separation progressive framework, NAMI. We divide the rectified flow into different stages based on resolution, connecting stages with BridgeFlow. Fewer transformer layers are used at the low-resolution stages to generate image layouts and concept contours, while more layers are progressively added as the resolution increases. As shown in the Figure 2, we visualize the image generation process of FLUX-dev [21] and NAMI-2B under the same inference steps. Despite using a lower resolution and fewer model parameters in the early stages, NAMI-2B still effectively generates image layouts and concepts. Furthermore, considering the overall process, NAMI-2B achieves a faster generation speed.
Current benchmarks such as GenEval [11], DPG-Benchmark [18], and ImageReward [51], are primarily used to evaluate a model’s performance. However, these benchmarks suffer from issues such as limited caption quantity, distribution biases. Therefore, a more comprehensive benchmark is essential for accurately evaluating a model’s capabilities. We combine open benchmarks, AI-generated, and human-created test prompts to construct a multidimensional evaluation dataset named NAMI-1K.
@@
    The NAMI-2B employs temporal separation and spatial cascading within the model to efficiently generate images. As shown in Figure 1, using the same model size, our approach reduces the time required to generate 1024 resolution images by 64%, while still maintaining a high level of image quality.
@@
    Furthermore, we propose a BridgeFlow module to align flows between different stages.
  4. 4.
@@ 2 Related Work
Diffusion models [44, 16, 38, 9] have become a powerful framework in image generation, with latent diffusion models further enhancing both quality and efficiency. DiT [31] has introduced a transformative change by replacing the conventional UNet [39, 41, 8] architecture used in models like SDXL [33] with a transformer [48]-based framework. This advancement has led to the development of models such as PixArt-$\alpha$ [5], Hunyuan-DiT [23], LUMINA-Next [55], SD3 [10], and FLUX [21], which efficiently integrate multimodal information to improve text-to-image generation, thereby demonstrating the effectiveness of transformer-based approaches. Notably, SD3 and FLUX employ rectified flow [27, 1, 24], which connects data and noise on a straight line, improving both training and inference efficiency. Additionally, the MM-DiT [10] block has been proposed, achieving state-of-the-art generation performance. Our work builds upon MM-DiT as the fundamental module while adopting rectified flow as a specific forward path selection strategy.
@@
Figure 3: Overview of NAMI: The left figure shows the progressive flow transformers of NAMI, where the same color represents the same module. The right figure depicts the integration of the BridgeFlow module, which establishes connections across adjacent time windows. Specifically, we divide the image generation process into $K$ resolution stages and the entire flow is divided into $K$ time windows, where adjacent stages are connected through upsampling and the BridgeFlow module. We use fewer transformer layers at the low-resolution stages to generate image layouts and concept contours, progressively adding more layers as the resolution increases.
@@ 3.1 Preliminary Study
$\min_{\theta}\mathbb{E}_{x_{0}\sim\mathcal{N}(0,I),x_{1}\sim D}\left[\int_{0}^{1}\left\|(x_{0}-x_{1})-v_{\theta}(x_{t},t)\right\|^{2}\,dt\right]$
@@ 3.2 Progressive Rectified Flow Transformers
Our method treats the image generation process as a transformation from a low-resolution image containing coarse-grained concepts to a high-resolution image with rich details. At low resolution, we employ piecewise flow [54] and partial modules to reduce computational costs. As shown in Figure 3 , we divide the image generation into $K$ resolution stages, where $Up(.,.)$ denotes the upsampling function and $Down(.,.)$ denotes the downsampling function. For flow processing, we refer to pyramid flow matching [19] and divide the entire flow into $K$ time windows $\{[t_{k-1},t_{k}]\}_{k=1}^{K}$, where $1=t_{K}>\dots>t_{k}>t_{k-1}>\dots>t_{0}=0$. Linear interpolation is also used between noise and data distribution, that is, $x_{t_{k}}=t_{k}x_{0}+(1-t_{k})x_{1}$. The starting point for each time window is given by $\hat{x}_{s_{k}}=BridgeFlow(Up(Down(x_{t_{k-1}},2^{k+1})))$, which is obtained by upsampling the endpoint of the previous stage. The end point is given by $\hat{x}_{e_{k}}=Down(x_{t_{k}},2^{k})$, which is obtained by downsampling $x_{1}$ and adding noise. The flow within the window is given by:
@@
$\min_{\theta_{k}}\sum_{k=1}^{K}E_{(k,t^{\prime},(\hat{x}_{s_{k}},\hat{x}_{e_{k}}))}\left[\int_{t_{k-1}}^{t_{k}}\left\|(\hat{x}_{s_{k}}-\hat{x}_{e_{k}})-v_{\theta_{k}}(\hat{x}_{t},t^{\prime})\right\|^{2}\,dt^{\prime}\right]$
3.3 BridgeFlow Module
To ensure the continuity of probabilistic paths across different stages, we design the BridgeFlow module with a linear transformation that aligns the probability distributions at stage boundaries. Specifically, for the endpoint of each stage, we first apply upsampling to match the resolution, followed by a linear transformation $\hat{x}_{s_{k}}=W\hat{x}_{e_{k-1}}+B$ to match the distribution of the subsequent stage’s starting point. In practice, every BridgeFlow module is pretrained with a mean squared error (MSE) loss according to the defined flow partitions.
@@ 3.4 Multi-Resolution Training
2: Create $K$ time windows $t_{k-1},t_{k}]$ for $k=1$ to $K$ with $t_{K}=1,t_{0}=0$
@@
5:  for each time window $t_{k-1},t_{k}]$ do
@@
9:   if $k=1$ then
10:    $\hat{x}_{s_{k}}\leftarrow x_{0}$
11:   else
12:    $\hat{x}_{s_{k}}\leftarrow\text{BridgeFlow}(\text{Up}(\text{Down}(x^{\prime}_{t_{k-1}},2)))$
13:   end if
14:   Compute the end point $\hat{x}_{e_{k}}=x^{\prime}_{t_{k}}$
15:   Sample $t^{\prime}\in[t_{k-1},t_{k}]$, compute $\hat{x}_{t}=t^{\prime}\hat{x}_{s_{k}}+(1-t^{\prime})\hat{x}_{e_{k}}$
16:   Compute the loss:
@@
17:  end for
18:  Weighting the losses for different time windows and perform backpropagation
19:until convergence
Considering knowledge sharing between modules and efficient training, we propose a multi-resolution progressive approach during the training process. Different resolutions of data can be used at different stages, and the optimization follows the approach in Equation 4. The training process is shown in Algorithm 1. Unlike the conventional method of training at low resolutions first and then fine-tuning at higher resolutions separately, our approach enables the joint training of data with different resolutions simultaneously. This facilitates knowledge sharing within the model and helps prevent catastrophic forgetting during fine-tuning at high resolutions. At each stage, we train using images with resolutions greater than or equal to the current resolution. The $Down(.,.)$ function is applied to downsample the images to the corresponding resolution, from which the starting and ending points are computed to obtain the target and loss for that stage. The loss of different stages is jointly optimized by dynamically adjusting their weights according to the training process. Additionally, during training, we also follow SD3 [10] and use logit-normal sampling for each stage. To enable CFG [15], we apply a 0.1 probability for random dropping of the prompt.
@@ 3.5 Inference
During inference, we sample a noise at the minimum resolution, start from stage $k=1$ and proceed until $k=K$, where at each stage, we use the Flow-Euler-discretization sampler. For the jump points between stages, We use upsampling and BridgeFlow to perform the transformation. Our approach can reduce 64% time during inference.
@@ 4.1 Model Details
4.2 Inference Time Details
As shown in Table 2, we provide the inference time of the NAMI-2B model at each resolution, using 10 inference steps per stage. The ”Overall” time additionally includes the upsampling and BridgeFlow modules at the connection points between stages. Compared to the Flux-based 2B model with the same total of 30 sampling steps, our inference time is significantly reduced 64%. The design of the flow piece according to resolution reduces the computation time by 53%, while the model partition further decreases the time by an additional 11%. Refer to Appendix C for details.
Method 256 512 1024 Overall Reduction
Baseline - - 8.47 8.47 -
NAMI-2B 0.27 0.45 2.21 2.98 64.82%
Table 2: The inference time details of the NAMI architecture (measured in seconds).
@@ 4.3 Experiments Details
We first pretrain the BridgeFlow modules between stages according to the time window division, using a learning rate of 1e-3. Convergence is typically reached within 10k training steps. Subsequently, we trained the NAMI-2B model using the multi-resolution simultaneous training strategy described in Method. At the early stage of training, we set the learning rate to 1e-4 and sampled three resolutions within each batch at a ratio of 4:2:1 (from low to high resolution) for 120k steps. Subsequently, we adjusted the sampling ratio to 2:4:1 and continued training for a further 80k steps. Finally, we modified the ratio to 1:2:4, reduced the learning rate to 5e-5, and trained for 50k steps to obtain the final NAMI-2B model.
@@ 4.3.1 Quantitative Comparison
Models(Params) Ovr Sgl Two Cnt Col Pos CA
FLUX-dev(12B) 0.67 0.99 0.81 0.79 0.74 0.20 0.47
FLUX-schnell(12B) 0.71 0.99 0.92 0.73 0.78 0.28 0.54
LUMINA-Next(2B) 0.46 0.92 0.46 0.48 0.70 0.09 0.13
SDXL(2.6B) 0.55 0.98 0.74 0.39 0.85 0.15 0.23
Hunyuan-DiT(1.5B) 0.63 0.97 0.77 0.71 0.88 0.13 0.30
SD3-medium(2B) 0.62 0.98 0.74 0.63 0.67 0.34 0.36
Sana(1.6B) 0.66 0.99 0.77 0.62 0.88 0.21 0.47
NAMI-2B(2B) 0.65 0.99 0.78 0.64 0.82 0.20 0.45
Table 3: Comparison of different methods on GenEval. With highlight the best, second best entries. Ovr & Sgl & Two & Cnt & Col & Pos & CA mean: Overall & Single & Two & Counting & Colors & Position & Color Attribution.
We found that GenEval [11] and DPG-Benchmark [18] primarily focus on text-image alignment, with limited prompt diversity.
Models(Params) Ovr Gbl Ent Attr Rel Oth
FLUX-dev(12B) 84.0 82.1 89.5 88.7 91.1 89.4
FLUX-schnell(12B) 84.8 91.2 91.3 89.7 86.5 87.0
LUMINA-Next(2B) 74.6 82.8 88.7 86.4 80.5 81.8
SDXL(2.6B) 74.7 83.3 82.4 80.9 86.8 80.4
Hunyuan-DiT(1.5B) 78.9 84.6 80.6 88.0 74.4 86.4
SD3-medium(2B) 84.1 87.9 91.0 88.8 80.7 88.7
Sana(1.6B) 84.8 86.0 91.5 88.9 91.9 90.7
NAMI-2B(2B) 83.8 90.3 89.4 88.3 88.0 90.6
Table 4: Comparison of different methods on DPG-Benchmark. With highlight the best, second best entries. Ovr & Gbl & Ent & Attr & Rel & Oth mean: Overall & Global & Entity & Attribute & Relation & Other
@@ 4.3.2 Human Evaluation
To address the limited prompt diversity of open-source benchmarks and to accurately evaluate the model’s performance, We propose the NAMI-1K evaluation dataset from the perspective of human preference performance.
Models(Params) Rele Cohe Aes Real Overall
Flux-dev(12B) 83.93 83.28 84.26 90.16 85.05
SD3-medium(2B) 75.74 65.90 61.64 75.74 69.97
Infinity(2B) 76.39 65.25 61.97 74.43 69.77
Hunyuan-DiT(1.5B) 74.43 69.51 63.93 64.92 68.95
SANA(1.6B) 75.41 62.30 60.00 72.46 67.80
NAMI-2B(2B) 76.07 66.89 62.30 76.72 70.69
Table 5: Human evaluation results on NAMI-1K dataset. Rele & Cohe & Aes & Real mean: Relevance & Coherence & Aesthetic & Realism. With highlight the best, second best entries.
@@
As shown in Figure 4, the prompt lengths in GenEval are primarily concentrated within 10 words, while those in the DPG-benchmark are mainly distributed between 50 and 80 words. In contrast, NAMI-1K exhibits a diverse distribution across different lengths within 120 words. As described in Appendix A, compared to GenEval and the DPG-benchmark, the topic distribution of NAMI-1K is more comprehensive and balanced.
@@
Evaluation Results The evaluation metrics are presented in Table 5. Although our model still trails behind FLUX-dev [21], which have considerably larger parameter sizes, it shows a clear advantage over SD3-medium [10], Infinity [13], Hunyuan-DiT [23], and SANA [50], which have comparable parameter sizes. As illustrated in Appendix D, a comparison of the generation results is provided between different methods.